By using the knowledge of relationships among the various costs of production concepts fill the blank space of the following table. 

Quantity TC TFC TVC ATC AVC MC

0 125

10 5

20 105

30 110

40 255

50 3

60 3


1
Expert's answer
2022-02-21T11:51:50-0500

Output 0 10 20 30 40 50 60 70

FC 125 125 125 125 125 125 125

"TFC=875"

VC 0 -120 -10 -15 130 -120 -122

"TVC=-257"

TC 125 5 105 110 255 3 3

AVC -0.5714 -0.0476 -0.074 0.6190 -0.5810 -0.5810

AFC 0.5952 0.5952 0.5952 0.5052 0.5952 0.5952

ATC 0.5952 0.0238 0.5476 0.5238 1.2142 0.0142 0,0142

MC _ -12 10 0.5 14.5 -25.2 -25.2

Where; "TC=FC+VC"

              "ATC=AVC+AFC"

"AVC=\frac{ variable cost}{TP}"

"AFC=\frac{fixed cost}{TP}"

"TP=210"